You received a year-end bonus of $17000 and invested it at 6.5%. How much additional money would you need to invest at 5% so that the return on the investments would be 6%
Answer by 24HoursTutor.com(40)

Let additional money to be invested be x
So, the equation becomes
17000*6.5% + x*5% = (17000+x)*6%
1105 + 0.05x = 1020 + 0.06x
0.01x = 85
x = 8500
The above question is answered by one of the experts from 24HoursTutor.com
We provide live online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr